\begin{thebibliography}{10}

\bibitem{1991-01a} 
Ashkenazi, J., Barnes, S.E., Zuo, F., Vezzoli, G.C., Klein, B.M.  Eds.
High-Temperature Superconductivity :Physical Properties, Microscopic Theory, and Mechanisms,
Proceedings  of the workshop  held  Jan.  3-9,  1991, University  of Miami, Coral  Gables, 
(Springer US, Boston, MA,1991)  doi:10.1007/978-1-4615-3338-2

\bibitem{1992-01a}
Bar-Yam, Y., Egami, T., Leon, J. M.-d., Bishop, A.R. (Eds.), 
Proc. of the conference: Lattice Effects in High-$T_c$  Superconductors. Santa Fe, New Mexico, January 13-15, 1992 
(World Scientific Pub., Singapore 1992) 

\bibitem{1992-05x} 
Muller, K.�A., Benedek, G.  (Eds.), Phase Separation in Cuprate Superconductors, 
Proc.  Erice Workshop, Italy 6-12 May 1992,
(World Scientific Pub., Singapore,1993)

\bibitem{1992-08x} 
Proc. 7th Int. Conf. X-ray Absorption Fine Structure, Kobe, Aug 1992.
Jpn. J. Appl. Phys.  {\bf 32} (Supplement 32-2) 1993
http://iopscience.iop.org/issue/1347-4065/32/S2

\bibitem{1994-10a} 
Sigmund, E.;  Muller, K.A. (Eds.),
"Phase Separation in Cuprate Superconductors : 
Proc. of the second international workshop on "Phase Separation in Cuprate Superconductors" September 4-10, 1993, Cottbus, Germany  
(Springer Berlin Heidelberg, 1994)

\bibitem{1993-12} 
Bianconi, A., Jul. 2001. 
Priority date (7 Dec 1993)
Process of increasing the critical temperature $T_c$ of a bulk superconductor 
by making metal heterostructures at the atomic limit. 
US Patent 6,265,019. 

\bibitem{1994-07} 
Bianconi, A., Jul. 1994. 
On the Fermi liquid coupled with a generalized Wigner polaronic CDW giving high Tc superconductivity.
 Solid State Communications  {\bf 91} (1), 1-5.
 http://dx.doi.org/10.1016/0038-1098(94)90831-1

\bibitem{ruani1995} 
Mihailovic, D., Muller K. A.,  Ruani G, (Eds.), 
 Anharmonic properties of high-$T_c$ cuprates 
(Proc. of the International Workshop on Anharmonic properties of High-Tc Cuprates, Bled, Slovenia, September 1-6, 1994)
(World Scientific, Singapore 1995)

\bibitem{tenyearsa}
Kaldis, E., Liarokapis, E., Muller, K. A. (Eds.), 
High-Tc Superconductivity 1996: Ten Years after the Discovery. Vol. 343 of NATO ASI Series. Springer Netherlands, 
(Proc. of the International Workshop on High-Tc Superconductivity 1996: Ten Years after the Discovery, Delphi, Greece, 1996) 
http://dx.doi.org/10.1007/978-94-011-5554-0\_16

\bibitem{tenyears}
Bianconi, A., Saini, N. L., Lanzara, A., Perali, A., Rossetti, T., Valletta, A., 1997. 
From a homogeneous $CuO_2$ plane to a superlattice of quantum stripes. 
In: \cite{tenyearsa} pp. 383-403.

\bibitem{stripes98x}
 Bianconi  A., Saini, N. L. (Eds.)
 Stripes and related phenomena
 Proceedings of the International Conference Stripes 98, June 4-7 1998,Rome, Italy,
(Kluwer Academics - Plenum Publisher 2000)
isbn:9780306464195

 \bibitem{superstripes1999c} 
Caprara, S., Sulpizi, M., Bianconi, A., Di Castro, C., Grilli, M., Jun. 1999. 
Single-particle properties of a model for coexisting charge and spin quasicritical fluctuations coupled to electrons. 
Physical Review B  {\bf 59}  (23), 14980-14991.
http://dx.doi.org/10.1103/physrevb.59.14980

 \bibitem{superstripes2000c} 
 Bianconi, A., Bianconi, G., Caprara, S., Di Castro, D., Oyanagi, H., Saini, N. L., Dec. 2000. 
 The stripe critical point for cuprates. 
 Journal of Physics: Condensed Matter  {\bf 12} (50), 10655-10666.
http://dx.doi.org/10.1088/0953-8984/12/50/326

\bibitem{superstripes} 
Bianconi, A., Aug. 2000. 
Superstripes. 
International Journal of Modern Physics B  {\bf 14}, (29-31), 3289-3297.
http://dx.doi.org/10.1142/S0217979200003769

\bibitem{superstripes2000a} 
Saini, N. L., Bianconi, A., Dec. 2000. 
Superstripes by anomalous X-Ray diffraction and angle resolved photoemission in BI2212. 
International Journal of Modern Physics B {\bf 14} (29-31), 3649-3655.
http://dx.doi.org/10.1142/S0217979200004179

\bibitem{superstripes2000b} 
Bianconi, A., Saini, N. L., Agrestini, S., Castro, D. D., Bianconi, G., Dec. 2000. 
The strain quantum critical point for superstripes in the phase diagram of all cuprate perovskites. 
Int. J. Mod. Phys. B  {\bf 14}1(29-31), 3342-3355.
 http://dx.doi.org/10.1142/s0217979200003812
 
 \bibitem{superstripes2000x} 
 Di Castro, D., Bianconi, G., Colapietro, M., Pifferi, A., Saini, N. L., Agrestini, 
 S., Bianconi, A., Dec. 2000. 
 Evidence for the strain critical point in high $T_c$ superconductors. 
 The European Physical Journal B  {\bf18} (4), 617-624.
 http://dx.doi.org/10.1007/s100510070010

\bibitem{superstripes2001} 
Bianconi, A., Di Castro, D., Bianconi, G., Saini, N. L., Feb. 2001. 
The strain quantum critical point for superstripes. 
AIP Conf. Proc.  {\bf 554} (1), 124-132.
 http://dx.doi.org/10.1063/1.1363067
 
 \bibitem{superstripes2001a} 
Agrestini, S., Saini, N. L., Bianconi, G., Bianconi, A., Sep. 2003. 
The strain of $CuO_2$ lattice: the second variable for the phase diagram of cuprate perovskites. 
Journal of Physics A: Mathematical and General  {\bf 36} (35), 9133-9142.
http://dx.doi.org/10.1088/0305-4470/36/35/302
 
 \bibitem{erice2003}
 Bianconi A. (Eds.) 2006,
 Symmetry and Heterogeneity in High Temperature Superconductors.
 (Proceedings of the NATO Advanced Study Research. Workshop, Erice,  Sicily, Italy October 4  10, 2003)
 NATO Science Series, Vol.  {\bf 214}. (Springer, Dordrecht, The Netherlands, 2006)
ISBN:978-1-4020-3987-4

\bibitem{2005x}
Muller,  K. A., Bussmann-Holder, A., 
Superconductivity in Complex Systems, 
in Structure and Bonding series, vol.  {\bf 114} (Springer Berlin Heidelberg, 2005)

 \bibitem{2009x}
Caivano, R., Fratini, M., Poccia, N., Ricci, A., Puri, A., Ren, Z.-A., Dong, 
X.-L., Yang, J., Lu, W., Zhao, Z.-X., Barba, L., Bianconi, A., Jan. 2009. 
Feshbach resonance and mesoscopic phase separation near a quantum critical 
point in multiband FeAs-based superconductors. 
Superconductor Science and Technology  {\bf 22} (1), 014004.
http://dx.doi.org/10.1088/0953-2048/22/1/014004
 
 \bibitem{2009y}
Ricci, A., Poccia, N., Ciasca, G., Fratini, M., Bianconi, A., Aug. 2009. 
The microstrain-doping phase diagram of the iron pnictides: Heterostructures at atomic limit. 
Journal of Superconductivity and Novel Magnetism  {\bf 22} (6), 589-593.
 http://dx.doi.org/10.1007/s10948-009-0473-x

  \bibitem{super2011}
 Bianconi, A., Apr. 2011. 
 Resonances and complexity: From stripes to superstripes. 
 Journal of Superconductivity and Novel Magnetism  {\bf 24} (3), 1117-1121.
http://dx.doi.org/10.1007/s10948-011-1142-4

  \bibitem{super2012}
Bianconi, A., Poccia, N., Jul. 2012. 
Superstripes and complexity in High-Temperature superconductors. 
Journal of Superconductivity and Novel Magnetism  {\bf 25} (5), 1403-1412.
 http://dx.doi.org/10.1007/s10948-012-1670-6

 \bibitem{super2013}
 Bianconi, A., Innocenti, D., Campi, G., Aug. 2013. 
 Superstripes and superconductivity in complex granular matter. 
 Journal of Superconductivity and Novel Magnetism  {\bf 26} (8), 2585-2588.
http://arxiv.org/abs/1304.6939

\bibitem{super2014}
Bianconi, A., 2014. 
Superstripes and percolating Nanoscale-Striped puddles in heterostructures at atomic limit. 
Journal of Superconductivity and Novel Magnetism  {\bf 27} (4), 909-912.
 http://dx.doi.org/10.1007/s10948-014-2516-1

\bibitem{super2015}
Bianconi, A., Apr. 2015. 
Superstripes in the low energy physics of complex quantum matter at the mesoscale.
Journal of Superconductivity and Novel Magnetism  {\bf 28} (4), 1227-1229.
http://arxiv.org/abs/1503.02966

 \bibitem{lif10} 
Bianconi,  A. 2013.
Quantum Materials: Shape Resonances in Superstripes, 
Nature Physics {\bf 9}, 536 

 \bibitem{leggett}
Leggett, A. J.,  Zhang, S., 2012
The BEC-BCS Crossover: Some History and Some General Observations.
In The BCS-BEC Crossover and the Unitary Fermi Gas, Vol. 836; Zwerger, W.,
Ed.; (Springer, Berlin Heidelberg, 2012).
doi:10.1007/978-3-642-21978-8\_2

\bibitem{1} 
Bednorz, J. G., Muller, K. A., Jun. 1986. 
Possible high $T_c$  superconductivity in the Ba-La-Cu-O system. 
Zeitschrift fur Physik B Condensed Matter  {\bf 64} (2), 189-193.  
http://dx.doi.org/10.1007/bf01303701

\bibitem{3} 
Wu, M. K., Ashburn, J. R., Torng, C. J., Hor, P. H., Meng, R. L., Gao, 
L., Huang, Z. J., Wang, Y. Q., Chu, C. W., Mar. 1987. 
Superconductivity at 93 K in a new mixed-phase Y-Ba-Cu-O compound system at ambient pressure. 
Physical Review Letters  {\bf 58} (9), 908-910.  
http://dx.doi.org/10.1103/physrevlett.58.908

\bibitem{buss}
Bussmann-Holder, A., Keller, H., Bianconi, A. (Eds.), 2017. 
High-$T_c$ Copper Oxide Superconductors and Related Novel Materials. 
Vol. 255. Springer International Publishing, Cham.
 http://dx.doi.org/10.1007/978-3-319-52675-1

\bibitem{chu}
Chu C. W. in Abbamonte A. et al., Superstripes 2017,  Bianconi A. Editor,
Science Series: Vol.  {\bf 11} (Superstripes Press, Rome, Italy)  Isbn: 9788866830696 
http://www.superstripes.net

\bibitem{5} 
Kresin, V. Z., Jun. 1987. 
On the critical temperature for any strength of the electron-phonon coupling. 
Physics Letters A  {\bf 122} (8), 434-438.
http://dx.doi.org/10.1016/0375-9601(87)90744-4

\bibitem{kres} 
Kresin, V., 2017. Paths to Room-Temperature superconductivity. 
J Supercond Nov Magn (2017)
http://doi.org/10.1007/s10948-017-4382-0

\bibitem{lifh3s-1} 
Bianconi, A., Jarlborg, T., 2015.
Lifshitz transitions and zero point lattice fluctuations in sulfur hydride showing near room temperature superconductivity
Novel Superconducting Materials  {\bf 1},  37 

\bibitem{lifh3s-2} 
Bianconi, A., Jarlborg, T., 2015
EPL (Europhysics Letters) {\bf 112}, 37001 
http://dx.doi.org/10.1209/0295-5075/112/37001

\bibitem{lifh3s-3} 
Jarlborg, T., Bianconi, A., 2016
Scientific Reports {\bf 6}, 24816  
http://dx.doi.org/10.1038/srep24816.

\bibitem{lifh3s-4} 
Bussmann-Holder, A., Koehler, J., Simon, A., Whangbo, M.-H., Bianconi, A., Perali, A., Jul. 2017. 
The road map toward Room-Temperature superconductivity.
Manipulating different pairing channels in systems composed of multiple electronic components. 
Condensed Matter  {\bf 2}(3), 24.   
http://dx.doi.org/10.3390/condmat2030024.

\bibitem{1981-02}
Chakraverty, B. K., 1981. 
Bipolarons and superconductivity. 
Journal de Physique 42 (9), 1351-1356.
http://dx.doi.org/10.1051/jphys:019810042090135100

\bibitem{ogg}
Ogg, R. A., Mar. 1946. 
Bose-Einstein condensation of trapped electron pairs. 
phase separation and superconductivity of Metal-Ammonia solutions. 
Physical Review 69 (5-6), 243-244.
http://dx.doi.org/10.1103/physrev.69.243

\bibitem{pauling}
Pauling, L., Apr. 1949. 
A Resonating-Valence-bond theory of metals and intermetallic compounds. 
Proceedings of the Royal Society A: Mathematical, Physical and Engineering Sciences 196 (1046), 343-362.
http://dx.doi.org/10.1098/rspa.1949.0032

\bibitem{1987-04} 
Bianconi, A., Congiucastellano, A. Desantis, M., Rudolf, P., Lagarde, P., Flank, A.M., Marcelli, A., Jun 1987
Cu $L_{2,3}$ XANES of the high $T_c$ superconductor $YBa_2Cu_3O_7$.
presented at the Symposium on High $T_c$  Superconductivity at the 7th General Conference 
of the Condensed Matter Division of the European Physical Society, Pisa, Italy, 8 Apr. 1987; 
$http://adsabs.harvard.edu/abs/1987eps..symp.....B$

\bibitem{1987-09a} 
Bianconi, A., CongiuCastellano, A., De Santis, M., Rudolf, P., Lagarde, P., Flank, A. M., Marcelli, A., Sep. 1987. 
$L_{2,3}$ XANES of the high $T_c$ superconductor $YBa_2Cu_3O_7$  with variable oxygen content. 
Solid State Communications  {\bf 63} (11), 1009-1013. 
http://dx.doi.org/10.1016/0038-1098(87)90650-8

\bibitem{1987-09c}
Bianconi, A., Castellano, De Santis, M., Politis, C., Marcelli, A., Mobilio, S., Savoia, A., Sep. 1987. 
Lack of delocalized Cu p states at the Fermi level in the high-$T_c$  superconductor $YBa_2Cu_3O_7$ 
by XANES spectroscopy. 
Zeitschrift fur Physik B Condensed Matter  {\bf 67} (3), 307-312. 
http://dx.doi.org/10.1007/bf01307254

\bibitem{1978}
Bianconi, A., Doniach, S., Lublin, D., 1978
X-ray Ca K edge of calcium adenosine triphosphate system and of simple Ca compunds
Chemical Physics Letters {\bf 59}, 121. 
http://dx.doi.org/10.1016/0009-2614(78)85629-2.

\bibitem{1979}
Bianconi, A., Jan. 1979. 
Core excitons and inner well resonances in surface soft x-ray absorption (SSXA) spectra. 
Surface Science  {\bf 89} (1-3), 41-50.

\bibitem{1980}
Bianconi, A., Nov. 1980.
Surface x-ray absorption spectroscopy: Surface EXAFS and surface XANES. 
Applications of Surface Science  {\bf 6} (3-4), 392-418. 
http://dx.doi.org/10.1016/0378-5963(80)90024-0

\bibitem{1986-03} 
J. Garcia, A. Bianconi, M. Benfatto, C. R. Natoli, 1986
Le Journal de Physique Colloques  {\bf 47} C8-49.
http://dx.doi.org/10.1051/jphyscol:1986807
 
\bibitem{1982-09}
Bianconi, A., Sep. 1982. 
Multiplet splitting of final-state configurations in x-ray-absorption spectrum of metal $VO_2$: 
Effect of core-hole-screening, electron correlation, and metal-insulator transition. 
Physical Review B  {\bf 26} (6), 2741-2747. 
http://dx.doi.org/10.1103/physrevb.26.2741

\bibitem{1985-02}
Marcelli, A., Bianconi, A., Davoli, I., Stizza, S., Feb. 1985. 
Localization mixing and / or hybridization in intermetallic compounds R$Pd_3$ (R = La, Ce, Pr, Nd, Sm) by XANES.
Journal of Magnetism and Magnetic Materials  {\bf 47-48}, 206-208. 
http://dx.doi.org/10.1016/0304-8853(85)90395-6

\bibitem{1986-02} 
Davoli, I., Marcelli, A., Bianconi, A., Tomellini, M., Fanfoni, M., Feb. 1986. 
Multielectron configurations in the x-ray-absorption near-edge structure of NiO at the oxygen K threshold. 
Physical Review B  {\bf 33} (4), 2979-2982.  
http://dx.doi.org/10.1103/physrevb.33.2979

\bibitem{1987-01}
Bianconi, A., Marcelli, A., Dexpert, H., Karnatak, R., Kotani, A., Jo, T., Petiau, J., Jan. 1987. 
Specific intermediate-valence state of insulating 4f compounds detected by $L_3$ x-ray absorption. 
Physical Review B  {\bf 35} (2), 806-812. 
http://dx.doi.org/10.1103/physrevb.35.806

\bibitem{1987-02}
Kotani, A., Okada, M., Jo, T., Bianconi, A., Marcelli, A., Parlebas, Feb. 1987. 
Many body effect in inner shell photoemission and photoabsorption spectra of La compounds. 
Journal of the Physical Society of Japan  {\bf 56} (2), 798-809.  
http://dx.doi.org/10.1143/jpsj.56.798

\bibitem{zaanen}
Zaanen, J., Sawatzky, G. A., Allen, J. W., Jul. 1985. 
Band gaps and electronic structure of transition-metal compounds. 
Physical Review Letters  {\bf 55} (4), 418-421.
http://dx.doi.org/10.1103/physrevlett.55.418

\bibitem{1987-09d}
Bianconi, A., Congiu-Castellano, A., De-Santis, M., Delogu, Gargano, Giorgi, Sep. 1987. 
Localization of Cu 3d levels in the high $T_c$  superconductor $YBa_2Cu_3O_7$ by Cu 2p x-ray photoelectron spectroscopy. 
Solid State Communications  {\bf 63} (12), 1135-1139.
http://dx.doi.org/10.1016/0038-1098(87)91063-5

\bibitem{1987-09e}
Fujimori, A., Takayama-Muromachi, E., Uchida, Y., Sep. 1987. 
Electronic structure of superconducting Cu oxides. 
Solid State Communications  {\bf 63}(9), 857-860.
 http://dx.doi.org/10.1016/0038-1098(87)90901-x

\bibitem{1988-02}
Bianconi, A., Budnick, J., Flank, A. M., Fontaine, A., Lagarde, P., Marcelli, 
A., Tolentino, H., Chamberland, B., Michel, C., Raveau, B., Demazeau, G., Feb. 1988.
Evidence of $3d^9$-ligand hole states in the superconductor
$La_{1.85}Sr_{0.15}CuO_4$ from $L_3$ x-ray absorption spectroscopy. 
Physics Letters A  {\bf 127} (5), 285-291.
http://dx.doi.org/10.1016/0375-9601(88)90698-6

\bibitem{1987-09b}
Bianconi, A., Clozza, A., Congiu-Castellano, A., Della-Longa, S., De-Santis, 
M., Di-Cicco, A., Garg, K., Delogu, P., Gargano, A., Giorgi, R., Lagarde, P., Flank, A. M., Marcelli, A., 1987. 
Experimental evidence of itinerant Cu($3d^9$)-Oxygen-hole 
many body configuration in the high-$T_c$  superconductor $YBa_2Cu_3O_7$.
(Proc. Special Adriatico Research Conference on High Temperature Superconductors 6-8 Jul 1987. Trieste, Italy) 
International Journal of Modern Physics B (IJMPB)  {\bf 1} (3-4), 853-862.
http://www.worldscinet.com/ijmpb/01/0103n04/S0217979287001213.html

\bibitem{1987-09f}
Bianconi, A., Clozza, A., Congiu Castellano, A., Della Longa, S., De Santis, M., Di Cicco, A., Garg, K., Delogu, P., Gargano, A., 
Giorgi, R., Lagarde, P., Flank, A. M., Marcelli, A., Dec. 1987.
Cu $3d^9$-ligand hole configuration in $YBa_2Cu_3O_7$ by x-ray spectroscopies.
Le Journal de Physique Colloques  {\bf 48}(C9), C9-1179 - C9-1184.
(Proc. of 14th International Conference on x-ray and inner-shell processes, Paris, September 14-18, 1987)
 http://dx.doi.org/10.1051/jphyscol:19879212
 
\bibitem{1987-10}
Bianconi A. in 
Kanamori, J., Kotani, A., editors 1988. 
Core-Level spectroscopy in condensed systems, 
Proceedings of the tenth Taniguchi international symposium, Kashikojima, Japan, October 19-23, 1987.
http://dx.doi.org/10.1007/978-3-642-83437-0

\bibitem{2} 
Bednorz, J. G., Muller, K. A., Jul. 1988. 
Perovskite-type oxides - The new approach to high- $T_c$  superconductivity. 
Reviews of Modern Physics  {\bf 60} (3), 585-600.
URL http://dx.doi.org/10.1103/revmodphys.60.585

\bibitem{emery} 
Emery, V. J., Jun. 1987. 
Theory of high-$T_c$ superconductivity in oxides. 
Physical Review Letters  {\bf 58} (26), 2794-2797.
http://dx.doi.org/10.1103/physrevlett.58.2794

\bibitem{2a} 
Muller, K. A., Dec. 1991.
The first five years of high-$T_c$ superconductivity. 
Physica C: Superconductivity  {\bf 185-189}, 3-10.
http://dx.doi.org/10.1016/0921-4534(91)91942-w
.
\bibitem{1988-06a}
Bianconi, A., Budnick, J., Chamberland, B., Clozza, A., Dartyge, E., Demazeau, G., De Santis, M., 
Flank, A. M., Fontaine, A., Jegoudez, J., Lagarde, P., Lynds, L. L., Michel, C., Otter, F. A., Tolentino, H., Raveau, B., Revcolevschi, A., Jun. 1988. 
$3d^9L$ states induced by doping in $La_{1.85}Sr_{0.15}CuO_4$ and in magnetic and non magnetic $M_1Ba_2Cu_3O_7$ (M = Gd, Ho and Y). 
(Proc. International Conference on High Temperature Superconductors and Materials and Mechanisms of Superconductivity, Interlaken, Switzerland, February 28 - March 4, 1988)
Physica C: Superconductivity  {\bf 153-155}, 113-114.
http://dx.doi.org/10.1016/0921-4534(88)90507-2

\bibitem{1988-06c}
Bianconi, A., De Santis, M., Di Cicco, A., Clozza, A., Congiu Castellano, A., Della Longa, S., Gargano, A., Delogu, P.,
Dikonimos Makris, T., Giorgi, R., Flank, A. M., Fontaine, A., Lagarde, P., Marcelli, A., Jun. 1988. 
Weight of $3d^9$ ligand hole configuration as function of oxygen content in $YBa_2Cu_3O_{6.5+x}$ 
by joint $L_3$ XAS and XPS. 
(Proc. International Conference on High Temperature Superconductors and Materials and Mechanisms of Superconductivity, Interlaken, Switzerland, February 28 - March 4, 1988)
Physica C: Superconductivity {\bf 153-155}, 115-116.
http://dx.doi.org/10.1016/0921-4534(88)90508-4

\bibitem{1988-06f}
Bianconi, A., Desantis, M., Flank, A., Fontaine, A., Lagarde, P., Marcelli, A., Katayamayoshida, H., Kotani, A., Jun. 1988.
 Determination of the symmetry of the $3d^9L$ states by polarized Cu $L_3$ XAS spectra of single crystal $YBa_2Cu_3O_{6.9}$. 
 Physica C: Superconductivity {\bf 153-155}, 1760-1761.
URL http://dx.doi.org/10.1016/0921-4534(88)90469-8

\bibitem{1988-10bm}
Bianconi, A., Marcelli, A. (Eds.), 1989. 
High $T_c$  superconductors electronic structure, 
Proceedings of the International Symposium on the Electronic Structure of High $T_c$  Superconductors, 
Rome, 5-7 October 1988. (Pergamon Press, Oxford 1989)
http://www.worldcat.org/isbn/9780080375427

\bibitem{1988-01x}
Sarma, D. D., Rao, C. N. R., Jan. 1988. 
Nature of the copper species in superconducting $YBa_2Cu_3O_7$. 
Solid State Communications {\bf 65} (1), 47-49.
 http://dx.doi.org/10.1016/0038-1098(88)90585-6

\bibitem{1988-04}
Nucker, N., Fink, J., Fuggle, J. C., Durham, P. J., Temmerman, W. M., Apr. 1988. 
Evidence for holes on oxygen sites in the high-$T_c$  superconductors $La_{2-x}Sr_xCuO_4$ and $YBa_2Cu_3O_{7-y}$
Physical Review B {\bf 37} (10), 5158-5163.
http://dx.doi.org/10.1103/physrevb.37.5158

\bibitem{1988-06d}
Sarma, D. D., Strebel, O., Simmons, C. T., Neukirch, U., Kaindl, G., Hoppe, R., Mïuller, H. P., Jun. 1988. 
Electronic structure of high-$T_c$  superconductors from soft-x-ray absorption. 
Physical Review B {\bf 37} (16), 9784-9787.
http://dx.doi.org/10.1103/physrevb.37.9784

\bibitem{1988-10a}
Balzarotti, A., De Crescenzi, M., Motta, N., Patella, F., Sgarlata, A., Oct. 1988. 
Energy loss study of the electronic structure of $YBa_2Cu_3O_7$ high $T_c$  superconductor. 
Solid State Communications {\bf 68} (4), 381-386.
http://dx.doi.org/10.1016/0038-1098(88)90299-2

\bibitem{1988-10b}
Balzarotti, A., De Crescenzi, M., Motta, N., Patella, F., Sgarlata, A., Oct. 1988. 
Valence charge fluctuations in $YBa_2Cu_3O_7$  from core-level spectroscopies. 
Physical Review B {\bf 38} (10), 6461-6469.
 http://dx.doi.org/10.1103/physrevb.38.6461

\bibitem{1989-01}
Fujimori, A., Jan. 1989. 
Character of doped oxygen holes in high-$T_c$  Cu oxides. 
Physical Review B {\bf 39} (1), 793-796.
http://dx.doi.org/10.1103/physrevb.39.793

\bibitem{1989-09}
de Groot, F. M. F., Grioni, M., Fuggle, J. C., Ghijsen, J., Sawatzky, G. A., Petersen, H., Sep. 1989. 
Oxygen 1s x-ray-absorption edges of transition-metal oxides. 
Physical Review B {\bf 40} (8), 5715-5723.
http://dx.doi.org/10.1103/physrevb.40.5715

\bibitem{1990-01}
Eskes, H., Tjeng, L. H., Sawatzky, G. A., Jan. 1990. 
Cluster-model calculation of the electronic structure of CuO: 
A model material for the high-$T_c$  superconductors. 
Physical Review B {\bf 41} (1), 288-299.
http://dx.doi.org/10.1103/physrevb.41.288

\bibitem{1990-07}
Bianconi A. in High Temperature Superconductivity, C. Ferdeghini, A. S. Siri editors, 
(Proceedings of the third Italian national meeting on High Temperature Superconductivity 
Genova 12-14 Feb 1990)
(World Scientific Publisher, Singapore, July 1990)
isbn, 9814611670, 9789814611671 

\bibitem{1991-03}
Ronay, M., Santoni, A., Schrott, A. G., Terminello, L. J., Kowalczyk, S. P., Himpsel, F. J., Mar. 1991. 
A new correlation for $T_c$  from Cu 2p absorption. Solid State Communications {\bf 77} (9), 699-702.
http://dx.doi.org/10.1016/0038-1098(91)90772-n
 
\bibitem{1991-11}
Bianconi, A., Della Longa, S., Li, C., Pompa, M., Congiu-Castellano, A., Udron, D., Flank, A. M., Lagarde, P., Nov. 1991. 
Linearly polarized Cu $L_3$-edge x-ray-absorption near-edge structure of $Bi_2CaSr_2Cu_2O_8$. 
Physical Review B {\bf 44} (18), 10126-10138.
http://dx.doi.org/10.1103/physrevb.44.10126

\bibitem{1991-12}
Pompa, M., Li, C., Bianconi, A., Congiu-Castellano, A., della Longa, S., Flank, A. M., Lagarde, P., Udron, D., Dec. 1991. 
Full multiple scattering analysis of linearly polarized Cu $L_3$-edge XANES of $La_2CuO_4$. 
Physica C: Superconductivity {\bf 184} (1-3), 51-64.
http://dx.doi.org/10.1016/0921-4534(91)91500-4
 
\bibitem{1992-04}
Chen, C. T., Tjeng, L. H., Kwo, J., Kao, H. L., Rudolf, P., Sette, F., Fleming, R. M., Apr. 1992. 
Out-of-plane orbital characters of intrinsic and doped holes in $La_{2-x}Sr_xCuO_4$. 
Physical Review Letters {\bf 68} (16), 2543-2546.
http://dx.doi.org/10.1103/physrevlett.68.2543

\bibitem{1992-08}
Bocquet, A. E., Mizokawa, T., Saitoh, T., Namatame, H., Fujimori, A., Aug. 1992. 
Electronic structure of 3d -transition-metal compounds by analysis of the 2 p core-level photoemission spectra. 
Physical Review B {\bf 46} (7), 3771-3784.
http://dx.doi.org/10.1103/physrevb.46.3771

\bibitem{1993-05}
van Veenendaal, M. A., Eskes, H., Sawatzky, G. A., May 1993. 
Strong nonlocal contributions to Cu 2p photoelectron spectroscopy. 
Physical Review B {\bf 47} (17), 11462-11469.
http://dx.doi.org/10.1103/physrevb.47.11462

\bibitem{1993-10}
Pellegrin, E., Nucker, N., Fink, J., Simmons, C. T., Kaindl, G., Bernhard, J., Renk, K. F., Kumm, G., Winzer, K., Oct. 1993.
Polarized x-ray-absorption study of $Tl_2Ba_2CaCu_2O_8$ and $Tl_2Ba_2Ca_2Cu_3O_10$. 
Physical Review B {\bf 48} (14), 10520-10523.
http://dx.doi.org/10.1103/physrevb.48.10520

\bibitem{1995-04}
Nucker, N., Pellegrin, E., Schwei, P., Sohmen, E., Fink, J., Molodtsov, S. L., Simmons, C. T., Domke, M., 
Kaindl, G., Frientrup, W., Chen, C. T., Erb, A., Mï½ller-Vogt, G., Apr. 1995.
 Site specific and doping dependent electronic structure of $YBa_2Cu_3O_x$ probed by O(1s) and Cu(2p) x-ray absorption spectroscopy. 
 Synthetic Metals {\bf 71} (1-3), 1563-1566.
 http://dx.doi.org/10.1016/0379-6779(94)02951-t
 
 \bibitem{1997-04}
 Merz, M., Nucker, N., Pellegrin, E., Schweiss, P., Schuppler, S., Kielwein, M., Knupfer, M., Golden,
 M. S., Fink, J., Chen, C. T., Chakarian, V., Idzerda, Y. U., Erb, A., Apr. 1997. 
 X-ray absorption spectroscopy of detwinned $Pr_xY_{1-x}Ba_2Cu_3O{7-y}$ single crystals:Electronic structure and hole distribution. 
 Physical Review B {\bf 55} (14), 9160.
 http://dx.doi.org/10.1103/physrevb.55.9160

\bibitem{2007-08}
Ghiringhelli, G., Brookes, N. B., Dallera, C., Tagliaferri, A., Braicovich, L., Aug. 2007. 
Sensitivity to hole doping of Cu $L_3$ resonant spectroscopies: Inelastic x-ray scattering and photoemission of $La_{2-x}Sr_xCuO_4$. 
Physical Review B {\bf 76} (8), 085116.
http://dx.doi.org/10.1103/physrevb.76.085116

\bibitem{2010-08}
Ugenti, S., Cini, M., Seibold, G., Lorenzana, J., Perfetto, E., Stefanucci, G., Aug. 2010. 
Particle-particle response function as a probe for electronic correlations in the p-d hubbard model. 
Physical Review B {\bf 82} (7), 075137.
 http://dx.doi.org/10.1103/physrevb.82.075137

\bibitem{2017-07}
Chainani, A., Sicot, M., Fagot-Revurat, Y., Vasseur, G., Granet, J., Kierren, B., Moreau, 
L., Oura, M., Yamamoto, A., Tokura, Y., Malterre, D., Jul. 2017. 
Evidence for weakly correlated oxygen holes in the highest-$T_c$
cuprate superconductor $HgBa_2Ca_2Cu_3O_8$. 
Physical Review Letters  {\bf 119} (5), 057001.
http://dx.doi.org/10.1103/physrevlett.119.057001

\bibitem{lif1} 
Valletta, A., Bardelloni, G., Brunelli, M., Lanzara, A., Bianconi, A., Saini, N. L., Aug. 1997.
$T_c$  amplication and pseudogap at a shape resonance 
in a superlattice of quantum stripes. 
Journal of Superconductivity {\bf10} (4), 383-387.
http://dx.doi.org/10.1007/bf02765723

\bibitem{lif2} 
Valletta, A., Bianconi, A., Perali, A., Saini, N. L., Dec. 1997.
 Electronic and superconducting properties of a superlattice of quantum stripes at the atomic limit. 
 Zeitschrift fur Physik B Condensed Matter {\bf 104} (4), 707-713.
 http://dx.doi.org/10.1007/s002570050513

\bibitem{lif3} 
Bianconi, A., Valletta, A., Perali, A., Saini, N. L., Feb. 1998. 
Superconductivity of a striped phase at the atomic limit. 
Physica C: Superconductivity  {\bf296} (3-4), 269-280. 
http://dx.doi.org/10.1016/s0921-4534(97)01825-x

\bibitem{lif4} 
Bianconi, A., Nov. 2005. 
Feshbach shape resonance in multiband superconductivity in heterostructures. 
Journal of Superconductivity {\bf 18} (5-6), 625-636.  http://dx.doi.org/10.1007/s10948-005-0047-5

\bibitem{lif5} 
Bianconi, A., Jan. 2006. 
Multiband superconductivity in high $T_c$  cuprates and diborides. 
Journal of Physics and Chemistry of Solids 
{\bf67} (1-3), 567-570. http://dx.doi.org/10.1016/j.jpcs.2005.10.160

\bibitem{lif6} 
Perali, A., Innocenti, D., Valletta, A., Bianconi, A., Sep. 2012.
Anomalous isotope effect near a 2.5 Lifshitz transition in a multi-band multi-condensate superconductor made of a superlattice of stripes.
Superconductor Science and Technology 
{\bf 25} (12), 124002.  http://arxiv.org/abs/1209.1528

\bibitem{lif8} 
Innocenti, D., Poccia, N., Ricci, A., Valletta, A., Caprara, S., Perali, A., Bianconi, A., Nov. 2010.
Resonant and crossover phenomena in a multiband superconductor: Tuning the chemical potential near a band edge. 
Physical Review B {\bf 82} (18), 184528.
http://dx.doi.org/10.1103/physrevb.82.184528

\bibitem{lif9} 
Innocenti, D., Caprara, S., Poccia, N., Ricci, A., Valletta, A., Bianconi, A., Jan. 2011.
Shape resonance for the anisotropic superconducting gaps near a Lifshitz transition: the effect of electron hopping between layers. 
Superconductor Science and Technology {\bf 24} (1), 015012. 
http://dx.doi.org/10.1088/0953-2048/24/1/015012

\bibitem{lif11}
Bianconi, A., Jul. 2013. 
Shape resonances in multi-condensate granular superconductors formed by networks of nanoscale-striped puddles. 
Journal of Physics: Conference Series {\bf 449} (1), 012002.
http://dx.doi.org/10.1088/1742-6596/449/1/012002

\bibitem{lifpter} 
Mazziotti, M. V., Valletta, A., Campi, G., Innocenti, D., Perali, A., Bianconi, A., May 2017. 
Possible Fano resonance for high $T_c$  multi-gap superconductivity in p-Terphenyl doped by k at the Lifshitz transition. 
EPL (Europhysics Letters) {\bf 118} (3), 37003.   http://dx.doi.org/10.1209/0295-5075/118/37003 

\bibitem{kus}  
Kusmartsev, F. V., Di Castro, D., BIanconi, G., Bianconi, A., 2000,
Physics Letters A  {\bf 275},118.
https://dx.doi.org/10.1016/s0375-9601(00)00555-7.

\bibitem{kug1} 
Kugel, K. I., Rakhmanov, A. L., Sboychakov,  A. O., Poccia, N., Bianconi, A., 2008
Phys. Rev. B. {\bf 78}, 165124.
https://dx.doi.org/10.1103/physrevb.78.165124.

\bibitem{kug2} 
 Kugel, K. I., Rakhmanov, A. L., Sboychakov, A.O., Kusmartsev, F.V., Poccia, N.,�Bianconi, A., 2009.
Superconductor Science and Technology {\bf 22}, 014007.
https://dx.doi.org/10.1088/0953-2048/22/1/014007 

\bibitem{kug3} 
Bianconi, A., Poccia, N., Sboychakov, A. O., Rakhmanov, A. L., Kugel, K., 2015.
Superconductor Science and Technology {\bf 28}, 024005,
https://dx.doi.org/10.1088/0953-2048/28/2/024005

\bibitem{jar}
Jarlborg, T., Bianconi, A., 2017.
 Electronic structure of $HgBa_2CuO_{4+\delta}$ with self-organized interstitial oxygen wires in the Hg spacer planes, 2017
J. Supercond. Nov. Magn. 
https://doi.org/10.1007/s10948-017-4453-2  (2017)

\bibitem{tall} 
Tallon, J. L., Talantsev, E. F., 2017. Compressed \& H$_3$S, superfluid density 
and the quest for Room-Temperature superconductivity. 
J. Supercond. Nov. Magn.  (2017).
http://doi.org/10.1007/s10948-017-4419-4

\bibitem{palu} 
 Palumbo F.,  Collective modes in BCS-like models, 2017, 
J. Supercond. Nov. Magn.
https://doi.org/10.1007/s10948-017-4420-y  (2017). 
 
 \bibitem{cha} 
Chavez, I., Garc�a, L. A., Grether, M., de Llano, M., Tolmachev, V. V., 2017. 
Extended BCS-Bose crossover, 
J. Supercond. Nov. Magn. https://doi.org/10.1007/s10948-017-4383-z (2017).
 
\bibitem{yana}
Yanagisawa, T., Miyazaki, M.,Yamaji, K.M., 
Crossover-induced spin fluctuation and electron pairing in strongly correlated electrons, 2017 
J. Supercond. Nov. Magn. (2017). https://doi.org/10.1007/s10948-017-4424-7

%TOPOLOGICAL AND SPIN-ORBITAL PHYSICS

\bibitem{ole} 
Brzezicki, W., Cuoco, M., Forte, F.,  Oles A. M., 
Topological Phases emerging from Spin-Orbital Physics,
 2017,
 J. Supercond. Nov. Magn.  (2017). https://doi.org/10.1007/s10948-017-4416-7

\bibitem{glo} 
Glodzik, S., G., Ptok, A., 2017. 
Quantum phase transition induced by magnetic impurity, 
J. Supercond. Nov. Magn.  https://doi.org/10.1007/s10948-017-4360-6 (2017). 

%CUPRATES

\bibitem{mar} 
Markiewicz, R. S., Buda, I. G., Mistark, P., Lane, C., Bansil, A., 2017. 
A new model of pseudogap physics in the cuprates. 
J. Supercond. Nov. Magn.  http://dx.doi.org/10.1007/s10948-017-4361-5 (2017).

\bibitem{tei} 
Gor'kov, L., Teitel'baum, G., 2017. On the origin of a small hole pocket in the Fermi surface of underdoped $YBa_2Cu_3O_y$.
J. Supercond. Nov. Magn.  (2017). http://doi.org/10.1007/s10948-017-4367-z

\bibitem{iva} 
Ivanov, A., Ivanov, V., Menushenkov, A., Wilhelm, F., Rogalev, A., Puri, A., Joseph, B., Xu, W., Marcelli, A., Bianconi, A., 
2017. Local noncentrosymmetric structure of $Bi_2Sr_2CaCu_2O_{8+y}$ 
by x-ray magnetic circular dichroism at Cu K-Edge XANES. 
J. Supercond. Nov. Magn. (2017). http://doi.org/10.1007/s10948-017-4418-5
                                                
\bibitem{mosk} 
Moskvin, A. S., Panov, Y., 2017. Topological structures in a model cuprate. 
J. Supercond. Nov. Magn. (2017). http://dx.doi.org/10.1007/s10948-017-4352-6

\bibitem{kan} 
Kanazawa, I., Maeda, R., 2017. 
Quantized massive gauge fields and anomalous angle-resolved photoemission spectra in High-$T_c$  cuprates. 
J. Supercond. Nov. Magn.  https://dx.doi.org/10.1007/s10948-017-4363-3 (2017). 

\bibitem{zhao} 
Zhao, H., Mou, Y., Feng, S., 2017. 
Correlation between charge order and second-neighbor hopping in cuprate superconductors. 
J. Supercond Nov. Magn. https://doi.org/10.1007/s10948-017-4327-7 (2017). 

%QUASI 1D SUPERCONDUCTORS

\bibitem{pto} 
Ptok, A., Cichy, A., Rodr�guez, K., Kapcia, K. J., 2017. 
Phase transitions in Quasi-One-dimensional system with unconventional superconductivity, 
J. Supercond. Nov. Magn. 
 https://doi.org//10.1007/s10948-017-4366-0 (2017). 

\bibitem{bar} 
Barba, L., Chita, G., Campi, G., Suber, L., Bauer, E., Marcelli, A., Bianconi, A., Oct. 2017. 
Anisotropic thermal expansion of p-Terphenyl: a Self-Assembled supramolecular array of poly-p-phenyl nanoribbons. 
J. Supercond. Nov. Magn. 
 https://dx.doi.org/10.1007/s10948-017-4407-8 (2017). 

%2D ELECTRON GAS 

\bibitem{ala} 
Alarco, J., Talbot, P., Mackinnon, I., 2017. 
A complete and accurate description of superconductivity of Al$B_2$-type structures from phonon dispersion calculations. 
J. Supercond. Nov. Magn. 
 http://doi.org/10.1007/s10948-017-4328-6 (2017).  

\bibitem{pud} 
Pudalov, V. M., Gershenson, M. E., 2017.
 Temperature dependence of renormalized spin susceptibility for interacting 2D electrons in silicon. 
J. Supercond. Nov. Magn.  https://doi.org/10.1007/s10948-017-4329-5 (2017). 

%IRON 

\bibitem{duan} 
Duan, C., Louca, D., 2017. 
Fe vacancy order and domain distribution in $A_xFe_{2-y}Se_2$, 
J. Supercond. Nov. Magn. 
http://doi.org/10.1007/s10948-017-4381-1  (2017).

\bibitem{umm} 
Ummarino, G. A., Daghero, D., Tortello, M., Gonnelli, R. S., 2017. 
Superconductivity on the verge of a Pressure-Induced Lifshitz transition in 
Ca$Fe_2$$As_2$: an interpretation within the Eliashberg theory. 
J. Supercond. Nov. Magn.   (2017). 
 http://doi.org/10.1007/s10948-017-4319-7

\bibitem{shy} 
Shylin, S., Ksenofontov, V., Naumov, P., Medvedev, S., Felser, C., 2017. 
Interplay between superconductivity and magnetism in Cu-Doped FeSe under pressure, 
J. Supercond. Nov. Magn.    https://doi.org/10.1007/s10948-017-4317-9  (2017). 

%OUT-OF-EQUILIBRIUM

\bibitem{tod} 
Toda, Y., Mochizuki, H., Tsuchiya, S., Kurosawa, T., Oda, M., Mertelj, T., Mihailovic, D., 2017. 
Nonequilibrium quasiparticle dynamics in Bi-Based superconductors measured by modulation photoexcitation spectroscopy. 
J. Supercond. Nov. Magn. https://doi.org/10.1007/s10948-017-4325-9 (2017). 

%SILVER 

\bibitem{groc} 
Grochala, W., 2017. Silverland: the realm of compounds of divalent silver and why they are interesting. 
J. Supercond. Nov. Magn.  https://doi.org/10.1007/s10948-017-4326-8 (2017). 

%GRANULar

\bibitem{mos} 
Moshe, A., Bachar, N., Lerer, S., Lereah, Y., Deutscher, G., 2017. 
Multi-Level Kondo effect and enhanced critical temperature in nanoscale granular Al. 
J. Supercond Nov. Magn.  https://doi.org/10.1007/s10948-017-4330-z (2017).

%ELECTRONIC SUPERCONDUCTING  QUANTUM DEVICES
\bibitem{sem} 
Semenov, A. G., Zaikin, A. D., 2017. Voltage noise in a superconducting wire with a constriction, 
J. Supercond. Nov. Magn.   https://doi.org/10.1007/s10948-017-4316-x (2017).

\bibitem{gala} 
Galaktionov, A., Golubev, D., Zaikin, A., 2017.
Intrinsic quantum dissipation in superconducting weak links. 
J. Supercond. Nov. Magn. https://doi.org/10.1007/s10948-017-4318-8 (2017)

\bibitem{marcelli} 
Marcelli, A., Coreno, M., Stredansky, M., Xu, W., Zou, C., Fan, L., Chu, W., Wei, S., Cossaro, A., Ricci, A., Bianconi, A., D Elia, A., Dec. 2017.
 Nanoscale phase separation and lattice complexity in $VO_2$:
 The Metal-Insulator Transition investigated by XANES via Auger electron yield at the vanadium $L_{2,3}$-edge and resonant photoemission.
 Condensed Matter {\bf 2} (4), 38.
 http://dx.doi.org/10.3390/condmat2040038
 
\end{thebibliography}

\documentclass[twocolumn,showpacs,preprintnumbers,amsmath,amssymb]{revtex4}
\usepackage{graphicx}
\usepackage{dcolumn}
\usepackage{bm}

\begin{document}

\title{ Lifshitz transitions in multi-band Hubbard models for topological superconductivity in complex quantum matter }
\author{Antonio Bianconi$^{1,2,3}$}

\affiliation{
$^1$RICMASS Rome International Center for Materials Science Superstripes, Via dei Sabelli 119A, 00185 Rome, Italy
\\
$^2$CNR-IC, Istituto di Cristallografia, Via Salaria Km 29.3, Monterotondo, Roma, I-00015, Italy
\\
$^3$National Research Nuclear University, MEPhI, Kashirskoe sh. 31, 115409 Moscow, Russia}

\date{31 Dec 2017, Journal of Superconductivity and Novel Magnetism DOI: 10.1007/s10948-017-4535-1}

\begin{abstract}

How the macroscopic quantum coherence can resist to the decoherence attacks 
of high temperature is a major challenge for the science of the 21st century. 
Superstripes 2017 conference held in Ischia on June 2017 has been focused on 
the new physics of high $T_c$  superconductors made of complex quantum matter.
Today the standard model of
high $T_c$ superconductivity which grabs the physics of complex quantum matter is
the multi-band Hubbard model where the dome of $T_c$ occurs 
by driving the chemical potential in the proximity of a topological Lifshitz transition.
The multi-gap superconductivity in the $T_c$ dome is driven by exchange interaction between
a first condensate in the BEC-BCS crossover which coexists with second BCS condensates.
The proximity to Lifshitz transitions in correlated electronic systems 
gives the ubiquitous arrested phase separation observed in all high temperature
superconductors. 
Non Euclidean filamentary hyperbolic geometry is needed for the space description of superstripes textures
produced by the coexistence of short range CDW puddles, 
hole poor SDW puddles and self organized dopants rich puddles. 
A road map to room temperature superconductors in particular organic compounds made of superlattices 
of quantum wires driven by Fano resonances with one of the condensates in the BEC-BCS crossover has been proposed. 
 
\end{abstract}

\pacs{05.70.Fh, 05.70.Jk,74.20.}

\maketitle
\section{Introduction.}

High temperature superconductivity has been found in a sequence of exotic complex materials: 
ceramics, inter-metallics, diborides, iron pnictides and chalcogenides, 
with the record for the highest critical temperature
$T_c$ in pressurized sulfur hydride near structural phase transitions.
These superconductors show a dome of high critical temperature
in a particular range of interstitials or defects concentration and in a particular range of pressure or misfit strain.
Each system shows a different complex landscape characterized by a different type
of multi-scale arrested phase separation with local-lattice-distortion, orbital, charge and spin modulations
forming textures of puddles of stripes from atomic-scale to nano-scale, mesoscale and micron-scale.
These exotic systems are clearly far away from a typical $conventional$ BCS superconductor 
made of homogeneous metal with a single large Fermi surface.

Experimental evidence for stripes due to anharmonic incommensurate lattice and orbital modulation has been 
found since 1990 by novel experimental methods using synchrotron radiation.
The early results have been reported and discussed 
at several international conferences focusing on lattice complexity
and phase separation \cite{1991-01a,1992-01a,1992-05x,1992-08x,1994-10a}.

The date of birth of the $stripes$ physics in high temperature superconductors can be fixed on Dec 7, 1993 
which is the priority date of the patent for material design
of heterostructures at atomic limit formed by superlattices 
of quantum stripes \cite{1993-12}. In this stripes scenario a Fermi liquid coexists with an incommensurate 1D 
charge density wave (CDW ) \cite{1994-07} forming a multigap superconductor near a Lifshitz transition where
the critical temperature amplification is driven by Fano resonances involving different condensates,
This stripes scenario for the high $T_c$ mechanism was presented at several international conferences in 1994-1996
\cite{ruani1995,tenyearsa}. 
The series of conferences on $stripes$ in high temperature superconductors started three years later on Dec 1996
 following the confirmation for the presence of stripes
by other standard experimental methods like neutron diffraction and 
NMR \cite{tenyears}, The first Stripes 1996 conference was 
followed by the second very large conference Stripes 1998 \cite{stripes98x} 
where the very simple Emery model of spin stripes 
with wave-vector $q_{sdw}$, locked with a charge stripes with wave-vector $q_{cdw}$= 2$q_{sdw}$ 
became very popular within the scientific community. 
The series of stripes conference have kept open the discussion on many different 
proposals for the stripes scenarios like 
the coexistence of short range charge stripes puddles unlocked from spin stripes puddles \cite{superstripes1999c}.
A major problem in the field was the diversity of the stripes scenarios in different families of hole doped cuprates 
which was solved in 2000 by the
disclosure of the key role of the lattice strain field with a critical strain value for the appearing of short range stripes ordering \cite{superstripes2000c}.
Moreover at the $stripes$ conference the term superstripes \cite{superstripes} has been coined to indicate 
the complex landscape generated by an arrested phase-separation near a critical strain point, 
forming a texture of  nano puddles of short range striped charge density wave order 
\cite{superstripes2000a,superstripes2000b,superstripes2000x,superstripes2001,superstripes2001a}.

 At that time, in the year 2000, the proposed heterogeneous landscape of superstripes was in contrast with the
 most popular paradigm  i.e., the single-band Hubbard model.
 Later many Scanning Tunneling Microscopy (STM) experiments have confirmed this scenario providing compelling 
 evidence for electronic nanoscale phase separation.
 Today the presence of nanoscale puddles of electronic pseudogap matter competing
 with superconducting condensate puddles with different symmetries \cite{erice2003} is well established.
In the new emerging paradigm high temperature superconductors 
 are described as particular cases of heterogeneous  complex quantum
 matter where particular forms of complexity are not detrimental (like normal disorder in the majority of disordered superconductors) but 
 they favor higher critical temperatures \cite{2005x}.

The name of the series of  $stripes$ conferences  changed its name into $superstripes$ conferences 
in 2008 driven by clear evidence 
for phase separation in iron based superconductors \cite{2009x,2009y}.
This decision marked the shift of the scientific interest toward 
quantum phenomena in complex matter \cite{super2011,super2012,super2013,super2014,super2015}.
and toward a new paradigm: the multigap superconductivity  near electronic 
topological Lifshitz transitions \cite{lif10} with pairing in one of the Fermi surface pockets is in the BEC-BCS crossover regime \cite{leggett} .

\section{From the  Woodstock of Physics to Superstripes 2017}

The international conference Superstripes 2017 has been held on June 4-10, 2017 at the Ischia island of the Neapolitan archipelago in Italy.
Scientists leaders in the field have been invited to discuss the latest advances in this field. 
Some discussions at Superstripes 2017 have given an answer to topics open since the 
APS March meeting held in New York, on 18 March 1987, 
the so called $Woodstock$ $of$ $Physics$, where
30 years ago Alex Muller presented the discovery
of high temperature superconductivity in ceramic La-Ba-Cu-O materials \cite{1}.
Paul C.W Chu reported superconductivity above liquid nitrogen temperature in $YBa_2Cu_3O_{6+y}$  ($Y123$) \cite{3}.
At Ischia 2017 conference Alex Muller, presented a review on the key role of complexity in cuprates recently  
reported in the book on high-$T_c$  copper oxide 
superconductors \cite{buss} presenting the scenario of high temperature superconductivity in complex materials.
Paul Chu discussed the role of interfaces in the enhancement of  $T_c$ above 77 K including the role of lattice architecture, 
internal strain and pressure. He pointed out the fact that a single-band Hubbard model 
is not sufficient for describing  high $T_c$,  why multi-band Hubbard models are needed \cite{chu}. 
Vladimir Kresin who proposed at the 1987 APS March 
meeting the theoretical scenario of strong  electron-phonon coupling in a complex lattice \cite{5} discussed in Ischia the  strong coupling 
limit in the pairing \cite{kres} involving high energy phonons
 in view of explaining 203 K superconductivity in pressurized $H_3S$ 
 which is today object of active research \cite{lifh3s-1,lifh3s-2,lifh3s-3,lifh3s-4}.

Band structure calculations of the parent compounds $La_2CuO_4$ and $YBa_2Cu_3O_6$ presented at the Woodstock of Physics 
predicted that the Fermi level is at half filling in a wide band due to 
the covalent bond between Cu(3d) [ml=2]  and O(2p$_{x,y}$) orbital in the metallic bcc $CuO_2$ layers.
Considering a perfect tetragonal bcc lattice
including only first-neighbor hopping t, the 2D Fermi surface at half filling is predicted
to have a square shape. The chemical potential is tuned
at the electronic topological Lifshitz transition
from the hole-like to the electron-like Fermi surface.
Here the system shows a peak in the Density of States (DOS) and strong electron-phonon scattering 
at the nesting wave-vector $2k_F$ connecting opposite sides of the square Fermi surface.
Therefore  high temperature superconductivity  was expected according with BCS theory
because of strong electron-phonon coupling and high density of states (DOS) at the van Hove singularity. 

On the contrary experiments have shown the failure of these predictions since
the parent compounds are antiferromagnetic insulators with an energy gap of about 2 eV. 
The insulating phase was explained by two different schools. 
The first school assumed the opening of 
a Peierls gap over the $full$ Fermi surface with wave-vector $2k_F$ 
associated with the formation of a 2D Peierls charge density wave.
The 2D Peierls CDW can be described as the ordering of polarons in the real space.
The CDW competes with 
the ordering of polaron pairs (bipolarons) in the k-space 
forming a superconducting phase in the strong coupling limit where below $T_c$ a Bose Einstein 
Condensation (BEC) occurs \cite{1981-02}.
The second  school proposed that the insulating phase was due to
the opening of a Hubbard gap $U_{dd}$ 
between [$3d^9$,$3d^9$]  and [$3d^8$,$3d^{10}$] configurations 
with the formation of a Mott insulator while the metallic superconducting phase 
was assigned to itinerant electronic singlets in a single band Hubbard model 
which condense like the preformed electron pairs in metal-ammonia solutions
 \cite{ogg} or like the resonating valence bond proposed by Pauling 
\cite{pauling} for transition metals.

\section{From single-band to multi-band Hubbard model and Lifshitz transitions}

It was well known that in the frame of a single-band Hubbard model the Mott insulator occurs if $U_{dd}$ is 
larger than the conduction band-width W. At the 1987 APS March meeting 
it was assumed by the scientific community that chemical doping form itinerant $Cu^{3+}$ impurity states,
with $3d^8$ configuration, moving in a background made of $Cu^{2+}$ ions with the $3d^9$ configuration.

Three weeks later on Apr 8th 1987 in the Symposium on High $T_c$ Superconductivity at the 7th General Conference 
of the Condensed Matter Division of the European Physical Society held in Pisa, Italy,
 it was reported \cite{1987-04} that the doped holes in doped 
 metallic $Y123$ do not form the expected $Cu^{3+}$ with
 $3d^8$ configuration but they form the 
 unexpected $Cu^{2+}$ $O^{1-}$ states called $3d^9L$, 
where L indicates the  hole in the ligand oxygen 2p orbital.

This result was obtained by Cu $L_{2,3}$-edge  x-ray absorption near edge structure (XANES) spectroscopy of the
high $T_c$ superconductor Y123  measured using ACO storage ring in Orsay, France in March 1987 \cite{1987-04,1987-09a} and 
by the Cu $K$-edge  XANES measured at the Adone storage
 ring in Frascati, Italy \cite{1987-09c}.

These results have been possible thanks to the development of XANES spectroscopy as a probe of 
both of multiple scattering resonances or shape resonances \cite{1978,1979,1980,1986-03}
and of many body electronic 
configurations  in valence 
fluctuation materials, heavy fermions and charge transfer correlated transition 
metal oxides like NiO, $CeO_2$ and $PrO_2$ \cite{1982-09,1985-02,1986-02,1987-01,1987-02,zaanen}.
                                                   
The  Cu $L_{2,3}$ X-ray photoelectron spectroscopy (XPS) spectra of Y123 were measured to get $U_{dd}$ in the Italian ENEA
Laboratory  \cite{1987-09d} and the results were found to be in agreement with the XPS experiments made by Fujimori in Tokyo \cite{1987-09e} 
at the same time.

These works have been confirmed in different cuprate families like in $La_{1.85}Sr_{0.15}CuO_4$ (La124)\cite{1988-02}.
The results have been presented at the four major international conferences in 1987:
1) the Pisa EPS April meeting \cite{1987-04}; 2) at the  Special Adriatico
Research Conference on High Temperature Superconductors held on 6-8 Jul 1987 in Trieste, Italy \cite{1987-09b}; 
3) at  the tenth Taniguchi international symposium held on October 19-23, 1987
in Kashikojima, Japan \cite{1987-09f};
and 4) at the 14th International Conference on x-ray 
and inner-shell processes, held in Paris on September 14-18, 1987  \cite{1987-10}.
The relevance of these results was recognized on 8 Dec 1987 at the Nobel prize ceremony
where Alex Muller \cite{2} reported that :
 
 [early photoelectron core-level spectra (XPS and UPS) by Fujimori 
  et al.\cite{1987-09e} and Bianconi et al.\cite{1987-09d}  in $La_{1-x}Sr_xCu_2O_{4-y}$ and
  $YBa_2Cu_3O_7$ did not reveal a final state owing
   to a $Cu^{3+}$ $3d^8$ state]
   
 In fact the  XPS experiments \cite{1987-09c,1987-09d} have clearly shown
that the parent compounds are Mott insulators with the Coulomb repulsion 
between two holes in the Cu 3d orbitals $U{dd}$= 6 eV  \cite{1987-09c,1987-09d}, larger than 
the conduction band width which was calculated by band structure calculations for non interacting fermions.
The Cu $L_3$ XANES experiments have shown that the carriers, created by doping,
are states with $3d^9 L$ many body configuration in the correlation gap.

The correlation gap in both La124 and Y123 compounds is not $U_{dd}$ as expected 
 for a single-band Hubbard model but the charge transfer gap for the excitation
 from the [$Cu(3d^9)$,$O(2p^6)$] to the  [$Cu(3d^{10})$,$O(2p^5)$] many 
 body configuration, called also the $3d^9$ to $3d^{10}L$ gap, which was predicted for the correlated charge transfer transition metal oxides
 \cite{zaanen}. 
 
A month after the EPS 1987 Pisa conference, Emery understood that the results of the 
Cu $L_3$-edge experiment  \cite{1987-04}
had falsified the single-band Hubbard model and proposed 
 in June 1987 the three-band Hubband model \cite{emery} involving p and d orbitals
for hole doped cuprates.

The experimental evidence of the  $3d^9L$ by A. Bianconi \cite{1987-09b} and  the V. Emery theory paradigm 
of the p-d three-band Hubbard model were presented together at the Adriatico Trieste meeting in July 1987.
 After the Adriatico 1987 meeting the fact that doped holes are in the oxygen orbital in cuprates was 
widely accepted by the community as recognized by Alex Muller in the opening talk at the 3rd international conference on 
Materials and Mechanisms of Superconductivity in 1991 at Kanazawa Japan \cite{2a} :

[The carriers in these type of conducting cuprates are holes on the oxygens. 
To me the first indication came from X-ray absorption spectroscopy by Prof. Bianconi at University of Rome] 

 Evidence for the $3d^9L$ states induced by doping in the correlation gap was 
presented in 1988 at the  first  $M^2S$ conference in Interlaken \cite{1988-06a,1988-06c,1988-06f}.
and at the international Symposium on the Electronic Structure of High $T_c$  Superconductors, 
Rome, 5-7 October 1988 \cite{1988-10bm}, 
where it was confirmed by many authors \cite{1988-01x,1988-04,1988-06d,1988-10a,1988-10b}. 
These results have been conformed by experiments and theories published
in 1989 \cite{1989-01,1989-09}; 
in 1990 \cite{1990-01,1990-07}; in 1991 \cite{1991-03,1991-11,1991-12}; 
in 1992  \cite{1992-04,1992-08}; 
in 1993 \cite{1993-05,1993-10}; 
in 1995-1997 \cite{1995-04,1997-04} 
and in these last 10 years \cite{2007-08,2010-08,2017-07}.

It was established that the correct model for high temperature superconductors 
is the multi-band Hubbard  model, where the domes of high $T_c$ 
occur in the proximity of electronic topological transitions called Lifshitz transitions.

Boguliobov equations for a multigap superconductor near a band edge have been solved by numerical calculations 
for the cases of sulfur hydrides, cuprates, diborides, and iron based superconductors \cite{lifh3s-1,lif1,lif2,lif3,lif4,lif5,lif6,lif8,lif9,lif11}. 
Recently it was shown that this mechanism for high $T_c$ can be in action in particular  filamentary organic  
compounds \cite{lifpter} where the condensate in the new appearing Fermi surface is in the BEC-BCS crossover regime.

Also the ubiquitous nanoscale phase separation  occurring 
in high temperature superconductors has been shown to be driven
 by the proximity to a Lifshitz transition in multi-band Hubbard models \cite{kus,kug1,kug2,kug3}.

\section{Recent advances}

At superstripes 2017 conference it was proposed that the origin of 
the multiple Fermi surfaces could be determined by oxygen interstitials self-organization 
as it has been shown  in the case of HgBa$_2$CuO$_{4+y}$ (Hg1201)  where the oxygen interstitials (O-i)  
are not homogeneously distributed but form one-dimensional atomic wires which can dramatically enhance the critical temperature. \cite{jar}.
The complex interplay of charge, spin \cite{tall,palu,cha,yana} and orbital degrees
 of freedom including spin-orbit interaction in topological matter \cite{ole,glo} 
 are now object of investigation as key ingredients. 
 The rich physics of unconventional superconductivity in cuprates and
their asymmetry between electron doped and hole doped families continue to attract theory and experimental research
\cite{mar,tei,iva,mosk,kan,zhao}. Ivanov has shown the noncentrosymmetric structure of a Bi2212 crystal at optimum doping
 in a high magnetic field by using x-ray magnetic circular dichroism, XMCD.
This new result suggests a possible role of spin-orbit coupling. 
Investigation of  basic physics in low dimensional quasi 1D 
superconductors \cite{pto,bar} and in the 2D electron gas  
\cite{ala,pud} have been object of high interest to unveil key features of dimemsionality in unconventional superconductors. 
Iron based superconductors continue to provide a clear 
case for the study of the interplay of nanoscale phase separation and multi-gap superconductivity \cite{duan,umm,shy}. 
New topics in physics and materials science are provided 
by systems out-of-equilibrium \cite{tod}, novel silver based materials \cite{groc} and granular materials \cite{mos}. 
Finally key advances in the quantum electronics using superconducting devices \cite{sem,gala} and on metal-to-insulator transition in $VO_2$ \cite{marcelli} have been reported at Superstripes 2017 

\section{Conclusions}

There is today after many years of discussion a growing agreement 
in the scientific community on some key common physical features of high temperature superconductors. 
All  high $T_c$ superconductors are highly inhomogeneous and different
multi-band Hubbard models are needed to describe unconventional high temperature superconductors.
Topology, Lifshitz transitions, and spin-orbit coupling are opening new fields of research in condensed matter
 of strongly correlated materials. Further works will be addressed to exchange interaction giving Fano
  resonances between condensates in the strong coupling regime in hot spots of the Fermi surface 
  coexisting with condensates in other portions of the Fermi surfaces in the weak coupling regime, which can be manipulated in 
  metalorganic materials to get room temperature superconductors \cite{lifpter}.

\end{document}